# Aberdeen Proving Ground, USA

### v4

@@ 4.2.2 Geometry of Pareto optimality
$\begin{split}&d_{avg}(f,g)\approx\sqrt{(r_{1}(C_{r_{1}})+\Delta_{r_{1}}/2)^{2}+(r_{2}(C_{r_{2}})+\Delta_{r_{2}}/2)^{2}}\\ &+[\Delta_{r_{2}}^{3}\sinh^{-1}(\Delta_{r_{1}}/\Delta_{r_{2}})+\Delta_{r_{1}}^{3}\sinh^{-1}(\Delta_{r_{2}}/\Delta_{r_{1}})\\ &+2\Delta_{r_{1}}\Delta_{r_{2}}\sqrt{\Delta_{r_{1}}^{2}+\Delta_{r_{2}}^{2}}]/(12\Delta_{r_{1}}\Delta_{r_{2}}),\end{split}$ (31)
@@
$\Delta_{r_{1}}=r_{1}(C_{r_{2}})-r_{1}(C_{r_{1}})$ (32)
and
$\Delta_{r_{2}}=r_{2}(C_{r_{1}})-r_{2}(C_{r_{2}}).$ (33)
@@
The average deviation from conjugacy over all points in the rectangle depicted in Fig. 1 is given by the integral
[…]
The integral can be split into the sum of two parts $d_{avg}=d_{avg,1}+d_{avg,2}$: the first part is given by the distance from the origin to the center of the rectangle depicted in Fig. 1, and second is the average distance from all points in the rectangle to the rectangle’s center. The first part, denoted by $d_{avg,1}$ follows directly from the geometry of Fig. 1 and is given by
$d_{avg,1}=\sqrt{(r_{1}(C_{r_{1}})+\Delta_{r_{1}}/2)^{2}+(r_{2}(C_{r_{2}})+\Delta_{r_{2}}/2)^{2}},$ (37)
where $\Delta_{r_{1}}$ and $\Delta_{r_{2}}$ are the lengths of the rectangle’s sides.
The second part $d_{avg,2}$ follows from rectangle_dist and is given by
$\begin{split}&d_{avg,2}=[\Delta_{r_{2}}^{3}\sinh^{-1}(\Delta_{r_{1}}/\Delta_{r_{2}})+\Delta_{r_{1}}^{3}\sinh^{-1}(\Delta_{r_{2}}/\Delta_{r_{1}})\\ &+2\Delta_{r_{1}}\Delta_{r_{2}}\sqrt{\Delta_{r_{1}}^{2}+\Delta_{r_{2}}^{2}}]/(12\Delta_{r_{1}}\Delta_{r_{2}})\end{split}$ (38)
@@ 4.4 Solutions in $\Psi$-space
While the primary focus here is computing deviation from conjugacy pseudometrics, one may be interested in analyzing systems in $\Psi$-space. Therefore, for the sake of completeness transformations in $\Psi$-space are recovered by substituting Eqs. 40 and 44 into Eq. 14. Since solutions for $T$ are based on left-eigenvectors which are not unique, an optimal magnitude for the eigenvectors must be found. Recognizing that $W_{g}$ can be arbitrarily scaled by a diagonal matrix $\Omega$, Eq. 14 becomes
@@ 5.1 Benchmarking example
Consider the systems $f$ and $g$ given by Eqs. 49 and 50 respectively.

### v6

@@ 4.2.2 Geometry of Pareto optimality
$\begin{split}&d_{avg}(f,g)\approx[r_{1}(C_{r_{2}})r_{2}(C_{r_{1}})A(2r_{1}(C_{r_{2}}),2r_{2}(C_{r_{1}}))-\\ &r_{2}(C_{r_{2}})r_{1}(C_{r_{2}})A(2r_{2}(C_{r_{2}}),2r_{1}(C_{r_{2}}))-r_{2}(C_{r_{1}})r_{1}(C_{r_{1}})A(2r_{2}(C_{r_{1}}),2r_{1}(C_{r_{1}}))\\ &+r_{1}(C_{r_{1}})r_{2}(C_{r_{2}})A(2r_{1}(C_{r_{1}}),2r_{2}(C_{r_{2}}))]\\ &/[r_{1}(C_{r_{2}})r_{2}(C_{r_{1}})-r_{2}(C_{r_{2}})r_{1}(C_{r_{2}})-\\ &r_{2}(C_{r_{1}})r_{1}(C_{r_{1}})+r_{1}(C_{r_{1}})r_{2}(C_{r_{2}})]\end{split}$ (31)
@@
$\begin{split}&A(\Delta_{1},\Delta_{2})=\Delta_{2}^{3}\sinh^{-1}(\Delta_{1}/\Delta_{2})+\Delta_{1}^{3}\sinh^{-1}(\Delta_{2}/\Delta_{1})\\ &+2\Delta_{1}\Delta_{2}\sqrt{\Delta_{1}^{2}+\Delta_{2}^{2}}]/(12\Delta_{1}\Delta_{2}).\end{split}$ (32)
@@
The average deviation from conjugacy over all points in the rectangle depicted in Fig. 1 is given by Eq. (31) which follows directly from rectangle_dist .
@@ 4.4 Solutions in $\Psi$-space
While the primary focus here is computing deviation from conjugacy pseudometrics, one may be interested in analyzing systems in $\Psi$-space. Therefore, for the sake of completeness transformations in $\Psi$-space are recovered by substituting Eqs. 36 and 40 into Eq. 14. Since solutions for $T$ are based on left-eigenvectors which are not unique, an optimal magnitude for the eigenvectors must be found. Recognizing that $W_{g}$ can be arbitrarily scaled by a diagonal matrix $\Omega$, Eq. 14 becomes
@@ 5.1 Benchmarking example
Consider the systems $f$ and $g$ given by Eqs. 45 and 46 respectively.